### v2

@@ 2.1.1 Backbone: Dual Parallel Branches
The design rationale for maintaining parallel branches rather than early fusion stems from the distinct noise characteristics and information density profiles of the two modalities. Depth measurements from structured-light or time-of-flight sensors exhibit modality-specific artifacts including edge bleeding, missing values at specular surfaces, and quantization noise at extended ranges—phenomena that bear no correspondence to the noise distributions present in RGB imagery. By preserving independent encoding pathways through the $P_{5}$ scale, the network permits each branch to develop specialized denoising and feature extraction strategies tailored to its respective input modality.
At the $P_{3}$ and $P_{4}$ scales within each branch, Semantic Recomposition Modules are inserted immediately following the C3k2 feature extraction blocks. These modules serve a critical preparatory function for subsequent cross-modal fusion by projecting branch-specific features into a normalized dictionary space where semantic concepts can be explicitly disentangled and refined. The placement at intermediate scales rather than solely at the deepest layer reflects the particular requirements of small object detection: $P_{3}$ and $P_{4}$ features retain sufficient spatial resolution to preserve fine-grained structural details essential for localizing compact defects, while possessing adequate semantic abstraction to support meaningful cross-modal correspondence. The branch-level SRM instances employ a compact dictionary configuration of 128 atoms, as the objective at this stage is intra-modal feature refinement rather than comprehensive cross-modal semantic modeling.
The terminal layer of each branch consists of a Spatial Pyramid Pooling-Fast module applied at the $P_{5}$ scale, which aggregates multi-scale contextual information through parallel max-pooling operations with kernel size 5. This configuration expands the effective receptive field without introducing additional downsampling, thereby enriching the semantic content of the deepest features while preserving the spatial correspondence necessary for accurate localization in subsequent detection stages.
2.1.2 Fusion Layers
The fusion architecture implements a hierarchical integration strategy that combines RGB and depth features at the $P_{4}$ and $P_{5}$ scales, where semantic abstraction is sufficiently mature to support meaningful cross-modal alignment while retaining adequate spatial granularity for precise defect localization. The decision to exclude $P_{3}$ from direct fusion and instead route it exclusively through the RGB pathway to the detection head reflects an empirical observation that shallow-scale features benefit more from modality-specific spatial detail preservation than from premature cross-modal mixing, which can introduce alignment noise that degrades small object detection performance.
At the $P_{4}$ fusion point, the network concatenates the SRM-processed features from both branches along the channel dimension, yielding a combined representation of doubled channel depth. This concatenated tensor is subsequently refined through a C3k2 block that learns to extract cross-modal complementary patterns—for instance, associating RGB texture signatures of corrosion or mechanical damage with corresponding depth discontinuities indicative of surface deformation. The C3k2 architecture is particularly suited to this fusion stage, as its cross-stage partial connectivity enables efficient gradient flow while the dual-path internal structure can implicitly develop specialized processing for RGB-derived and depth-derived feature subsets within the concatenated input.
Following initial C3k2 refinement, a second Semantic Recomposition Module is applied to the fused $P_{4}$ representation. Unlike the branch-level SRM instances that operate with 128-atom dictionaries for intra-modal purification, this fusion-stage module employs an expanded 256-atom dictionary to accommodate the richer semantic vocabulary required for modeling cross-modal correspondences. The dictionary learning mechanism enables explicit encoding of joint RGB-D semantic concepts that cannot be adequately represented within either individual modality—such as the characteristic combination of discoloration patterns in RGB imagery with geometric protrusions in depth maps that jointly indicate bird nest accumulation on transmission infrastructure.
The $P_{5}$ fusion pathway follows an analogous concatenation-refinement pattern, with the notable addition of a Contextual Semantic Integration Framework inserted between the C3k2 block and the terminal SRM. This module augments the fusion process with attention-based global context modeling, addressing the limitation that purely convolutional fusion operates within bounded receptive fields and may consequently fail to capture long-range dependencies essential for understanding defect patterns that span extended spatial regions. The attention mechanism computes pairwise affinities across spatial positions, enabling features at any location to directly aggregate information from semantically related positions throughout the feature map. This capability proves particularly valuable for transmission-line defect detection, where contextual relationships between components—such as the spatial configuration of insulators relative to conductors—provide critical diagnostic cues.
The progressive fusion strategy, wherein $P_{4}$ features are fused prior to $P_{5}$ and both fusion outputs are subsequently connected through the detection head’s feature pyramid, establishes information flow paths that enable deep-layer semantic guidance to inform shallow-layer feature refinement during backpropagation. This design ensures that the learning objective for small object detection at high-resolution scales can influence the cross-modal alignment learned at deeper fusion stages, thereby producing fused representations that are explicitly optimized for the downstream detection task rather than generic feature similarity.
2.1.3 Head
The detection head implements a bidirectional feature pyramid architecture that combines top-down Feature Pyramid Network pathways with bottom-up Path Aggregation Network connections, establishing comprehensive multi-scale feature communication essential for detecting transmission-line defects that manifest across a wide range of spatial extents. This architectural choice is motivated by the observation that transmission infrastructure inspection datasets exhibit extreme scale variation: defects range from millimeter-scale crack initiations visible only at high resolution to meter-scale structural deformations that require broad contextual understanding.
The FPN pathway initiates from the SRM-processed fused $P_{5}$ features, which are upsampled by factor 2 through nearest-neighbor interpolation and concatenated with the fused $P_{4}$ representation. A C3k2 block refines this combined tensor, learning to propagate high-level semantic information to the intermediate scale while integrating the finer spatial details preserved at $P_{4}$ resolution. This refined intermediate representation undergoes further upsampling and concatenation with the RGB branch’s $P_{3}$ features, producing the highest-resolution detection features at $P_{3}/8$ scale. The asymmetric design wherein $P_{3}$ receives input solely from the RGB branch rather than from a fused source reflects the observation that depth map resolution limitations and sensor noise characteristics render depth features less reliable at shallow scales, where RGB texture information provides superior discriminative power for small defect detection.
The subsequent PAN pathway reverses the information flow direction, propagating fine-grained spatial details from $P_{3}$ back toward deeper scales through strided convolutions and feature concatenation. This bottom-up augmentation addresses a fundamental limitation of purely top-down feature pyramids: while FPN effectively distributes semantic information to shallow layers, it cannot propagate the precise localization cues present at high resolution back to deeper layers that may benefit from such spatial detail during classification and regression. The PAN connections ensure that the $P_{4}$ and $P_{5}$ detection features receive both the semantic richness derived from deep fusion stages and the spatial precision necessary for accurate bounding box regression.
Each scale within the bidirectional pyramid terminates in a detection head that performs classification and localization for objects within a designated size range, with anchor-free prediction formulations that directly regress center coordinates, width, and height. The multi-head configuration enables scale-appropriate feature utilization: $P_{3}/8$ features detect small defects such as surface cracks and minor hardware damage, $P_{4}/16$ features handle medium-scale anomalies including insulator contamination and moderate structural displacement, while $P_{5}/32$ features address large defects encompassing conductor galloping damage and tower structural compromise. The architectural guarantee that each detection scale receives features informed by both cross-modal fusion semantics and multi-resolution spatial context establishes robust detection capability across the full spectrum of defect manifestations encountered in transmission-line inspection.
@@ 2.2 Semantic Recomposition Module
The Semantic Recomposition Module constitutes the foundational purification and enhancement mechanism deployed throughout the proposed network. As described in the architectural overview, SRM appears at multiple locations with differentiated configurations: compact-dictionary instances at branch $P_{3}$$P_{4}$ levels perform intra-modal semantic purification, while large-dictionary instances at fused $P_{4}$$P_{5}$ levels perform cross-modal semantic enhancement. This section elaborates the module’s internal design and explains how the unified encode-retrieve-normalize-decode mechanism serves these distinct functional roles.
SRM addresses a fundamental tension in multimodal feature processing: the need to suppress modality-specific noise while preserving discriminative details essential for detecting small-scale defects. Conventional approaches to feature refinement—such as channel attention or spatial attention mechanisms—operate by reweighting existing feature activations, which limits their ability to introduce genuinely new information into the representation or to enforce explicit constraints on what constitutes valid semantic content. In contrast, SRM adopts a dictionary-based retrieval paradigm that projects input features into a learned semantic codebook, performs structured refinement within that space, and injects the retrieved information back into the original representation through controlled residual mixing.
Figure 2: Architecture of the Semantic Recomposition Module. The input features are projected into a dictionary space via pointwise convolution, refined through depthwise spatial convolution, normalized at each spatial position (PONO denotes position-wise normalization), and decoded back to the original channel dimension. The output combines the refined features with the input through learnable residual mixing.
@@
The architecture of SRM, illustrated in Fig. 2, implements a four-stage pipeline followed by residual injection. Given an input feature map $\mathbf{F}\in\mathbb{R}^{B\times C\times H\times W}$, the encoding stage projects it into a lower-dimensional dictionary space through a $1\times 1$ pointwise convolution:
$\mathbf{E}=\mathcal{P}_{\mathrm{enc}}(\mathbf{F}),\quad\mathcal{P}_{\mathrm{enc}}:\mathbb{R}^{C}\rightarrow\mathbb{R}^{K},$ (1)
where $K$ denotes the dictionary size. This projection can be interpreted as computing soft assignment coefficients over a learned dictionary, with each spatial location in $\mathbf{E}$ encoding how strongly the corresponding input feature aligns with each of the $K$ dictionary atoms. The dictionary size $K$ is deliberately set smaller than the input channel count $C$ to encourage the network to learn a compact, semantically meaningful basis that captures essential structure while discarding irrelevant variations.
The retrieval stage aggregates spatial context within the dictionary space using a depthwise separable convolution with kernel size $5\times 5$:
$\mathbf{R}=\mathcal{D}_{5\times 5}(\mathbf{E}),$ (2)
where $\mathcal{D}_{5\times 5}$ operates independently on each dictionary channel without activation function. This design keeps the retrieval process close to an interpretable linear operation, aggregating information from local spatial neighborhoods for each atom while preventing nonlinearity from prematurely mixing noise with semantic information and spreading it across all channels. The enlarged kernel size ensures a receptive field sufficient to capture relationships between adjacent structural elements—such as the spatial arrangement of insulator components or the linear continuity of transmission conductors—while maintaining parameter efficiency through the depthwise formulation.
@@
The normalization stage applies position-wise standardization across the dictionary dimension:
@@
where $\epsilon$ ensures numerical stability. Unlike batch normalization or layer normalization, which compute statistics over spatial or batch dimensions, position-wise normalization treats each spatial location independently. This design choice is motivated by the observation that different regions of a transmission-line image exhibit vastly different statistical properties—background sky, metallic tower structures, and insulator assemblies each induce distinct activation patterns—and imposing shared statistics across these heterogeneous regions would suppress meaningful variations. For cross-modal scenarios, this normalization proves particularly crucial: RGB and depth modalities produce systematically different response magnitudes and local variances at corresponding spatial positions, and position-wise normalization explicitly aligns these distributions before the features proceed to subsequent processing, preventing the network from learning shortcuts based on statistical differences rather than semantic content.
The decoding stage maps the normalized dictionary representation back to the original channel dimension through another $1\times 1$ pointwise convolution:
@@
This projection reconstructs a refined feature map by linearly recombining the normalized dictionary atoms according to learned coefficients, translating the purified semantic information back into the representation space expected by subsequent network layers.
The final output combines the refined features with the original input through residual mixing:
@@
where the mixing coefficient $\alpha=0.8$ allows the dictionary-refined component to dominate while preserving a direct pathway for fine-grained information essential to small-object localization. Complete reliance on reconstruction risks boundary detail erosion, while excessive preservation of original features fails to suppress modality noise effectively; this convex combination strikes the necessary balance.
The differentiated deployment of SRM across the network exploits the same mechanism for distinct purposes through dictionary capacity variation. At branch $P_{3}$$P_{4}$ levels, the compact 128-atom dictionary creates an information bottleneck that forces features to be expressed using limited reusable primitives, effectively filtering out incidental noise and modality-private artifacts that cannot be stably reconstructed. This process creates a “reconstructable semantic subspace” for each modality, yielding features that are cleaner and more aligned in statistical properties before entering fusion. At fused $P_{4}$$P_{5}$ levels, the enlarged 256-atom dictionary accommodates the greater complexity of the shared feature space while maintaining reconstructability-based selection: shared semantic patterns that project cleanly into dictionary space and normalize consistently receive strong reinforcement, while modality-private residuals with inconsistent statistics receive attenuated injection. This mechanism transforms cross-modal alignment from implicit weight-based learning into explicit reconstruction-driven enhancement.
@@ 2.3 Contextual Semantic Integration Framework
While SRM effectively addresses intra-modal noise suppression and cross-modal distribution alignment through dictionary-based semantic purification, a fundamental limitation remains: the convolutional fusion operations at $P_{4}$ and $P_{5}$ levels operate within bounded receptive fields and consequently cannot capture long-range spatial dependencies. This limitation proves particularly problematic for transmission-line defect detection, where semantic interpretation frequently depends on global structural context. A damaged insulator, for instance, is more reliably identified when the network can reference the regular spatial arrangement of adjacent intact insulators to recognize deviation from expected patterns; similarly, foreign objects such as bird nests are distinguished from legitimate equipment by their inconsistency with the expected geometric layout of the transmission infrastructure. Without explicit mechanisms for global context modeling, the network must rely solely on local appearance cues that are often insufficient to disambiguate small defects from visually similar background elements.
To address this limitation, we introduce the Contextual Semantic Integration Framework (CSIF), deployed at the $P_{5}$ fusion stage where it serves to establish global contextual relationships among fused RGB-D features. The placement at $P_{5}$ reflects two design considerations: first, global attention is most valuable at deep semantic levels where features encode abstract concepts amenable to relational reasoning, whereas shallow features encoding fine-grained spatial details benefit more from local convolutional processing that preserves precise localization information; second, the computational cost of attention scales quadratically with spatial resolution, making the lowest-resolution $P_{5}$ scale the most tractable location for global modeling. However, directly applying standard Transformer-style attention to fused multimodal features introduces two subsidiary problems: full-channel global attention tends to cause detail erosion and over-smoothing that is particularly detrimental in small-object detection scenarios, and the concatenated RGB-depth features exhibit substantial distribution heterogeneity that standard layer normalization cannot adequately address. To resolve these issues, CSIF incorporates two specialized components: the Contextual Semantic Interaction Block (CSIB) implementing partial-channel attention with iterative refinement, and the Adaptive Semantic Regulation Mechanism (ASRM) providing structured normalization tailored for cross-modal feature distributions.
Figure 3: Architecture of the Contextual Semantic Integration Framework. The input features pass through a convolutional stem that harmonizes channel dimensions, followed by $n$ stacked Contextual Semantic Interaction Blocks that perform attention-based global modeling with interleaved normalization. An output projection aligns the refined features to the target dimension.
Fig. 3 illustrates the overall architecture of CSIF, which processes features through a “project–refine–project” pipeline designed to balance global context acquisition with detail preservation. Given an input $\mathbf{X}\in\mathbb{R}^{B\times C_{\mathrm{in}}\times H\times W}$ from the preceding fusion stage, CSIF produces a refined output $\mathbf{Z}\in\mathbb{R}^{B\times C_{\mathrm{out}}\times H\times W}$. The convolutional stem transforms the input into an internal embedding suitable for attention computation, expanding the channel dimension by factor $e$ to create sufficient capacity for multi-head attention while harmonizing the heterogeneous statistics of the incoming concatenated RGB-D features. To prevent the detail erosion problem inherent in full-channel attention, CSIF adopts a partial-channel attention design: the stem output is split into two portions along the channel dimension, with only one portion passing through the attention block sequence while the other maintains a stable bypass. This design constrains strong global interaction to a subspace, preserving a direct semantic pathway that retains fine-grained information essential for small-object localization while simultaneously stabilizing gradient flow during training. The core refinement consists of $n$ stacked CSIBs operating on the attention portion, with the iterative process formulated as $\mathbf{H}^{(i)}=\mathrm{CSIB}(\mathbf{H}^{(i-1)})$ for $i=1,\ldots,n$, where $\mathbf{H}^{(0)}$ denotes the initial attention portion. After $n$ iterations, the processed portion is concatenated with the bypassed portion, and an output projection maps the combined representation to the target channel dimension.
@@
Figure 4: Structure of the Contextual Semantic Interaction Block. Each block comprises a multi-head self-attention layer followed by ASRM normalization, then a feed-forward network followed by another ASRM normalization. Residual connections facilitate gradient flow at both stages.
Within the partial-channel attention framework, a subsidiary problem arises: how to effectively model global dependencies while maintaining stable feature distributions across the attention and feed-forward stages. Standard Transformer blocks employ pre-normalization, but this arrangement proves inadequate for cross-modal fusion scenarios where the concatenated RGB and depth features exhibit persistent distribution heterogeneity that channel-wise standardization alone cannot fully resolve. To address this problem, the Contextual Semantic Interaction Block adopts post-normalization placement with specialized ASRM modules that provide distribution-aware regularization. As shown in Fig. 4, each CSIB implements a two-phase cycle of attention and feed-forward transformation. The first phase applies multi-head self-attention to capture global spatial dependencies:
@@
where $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, $\mathbf{W}_{V}$ are learnable projection matrices and $d_{k}$ is the dimension per attention head. The number of heads adapts to the channel configuration as $h=\max(1,\lfloor C/64\rfloor)$ to ensure stable attention computation across varying feature dimensions. A residual connection preserves input information: $\mathbf{U}=\mathbf{H}+\mathrm{Attn}(\mathbf{H})$. Following attention, ASRM normalizes the feature distribution while preserving semantically relevant variations: $\hat{\mathbf{U}}=\mathrm{ASRM}_{1}(\mathbf{U})$. The second phase applies a feed-forward network for nonlinear feature transformation:
$\mathrm{FFN}(\hat{\mathbf{U}})=\mathbf{W}_{2}\cdot\sigma(\mathbf{W}_{1}\hat{\mathbf{U}}+\mathbf{b}_{1})+\mathbf{b}_{2},$ (9)
@@
The Adaptive Semantic Regulation Mechanism is introduced to stabilize feature distributions and regulate semantic magnitudes after global semantic interactions. In cross-modal fusion scenarios, global interactions such as self-attention and feed-forward transformations tend to amplify distributional discrepancies between heterogeneous semantic subspaces, which may result in feature drift and over-smoothing. ASRM addresses this issue by providing a controlled and learnable mechanism that balances distribution regularization with semantic preservation. Let the input feature be denoted as $\mathbf{x}\in\mathbb{R}^{B\times C\times H\times W}$. Rather than enforcing hard normalization, ASRM constructs a learnable semantic regulation gate that interpolates between the normalized representation and the original feature:
where $\mathrm{LN}(\cdot)$ denotes channel-wise layer normalization, $\boldsymbol{\gamma},\boldsymbol{\gamma}_{x}\in\mathbb{R}^{C\times 1\times 1}$ are learnable scaling parameters, and $\odot$ represents the Hadamard product. This formulation ensures that normalization contributes to representation learning in a regulated manner, instead of overriding the original semantic structure. On top of the regulated feature $\tilde{\mathbf{x}}$, ASRM applies a constrained semantic adaptation mapping $\mathcal{F}(\cdot)$, which operates in a low-dimensional semantic subspace and injects the adapted information back to the original feature space via a residual formulation:
@@
Through this residual semantic injection, ASRM confines global semantic corrections to a controlled subspace, preventing disruptive modifications to the backbone representation while still enabling effective regulation.
From an optimization perspective, ASRM exhibits progressive regulation behavior. By initializing $\boldsymbol{\gamma}$ to near-zero values and $\boldsymbol{\gamma}_{x}$ to unity, the mechanism approximates an identity mapping at early training stages, i.e., $\tilde{\mathbf{x}}\approx\mathbf{x}$, which avoids premature interference with unstable global attention patterns. As training proceeds, the learnable parameters adaptively increase the influence of the normalization branch, enabling a smooth transition from weak to strong semantic regulation. This progressive nature is particularly beneficial in cross-modal settings where the coupling between attention weights and normalization statistics requires gradual calibration. Conceptually, ASRM serves as a re-anchoring mechanism for global semantic updates: while self-attention captures long-range and cross-modal dependencies, ASRM constrains their impact through regulated normalization and residual semantic correction, ensuring that global context modeling enhances rather than obscures fine-grained discriminative cues. As a result, ASRM establishes a stable balance between global semantic coherence and local detail sensitivity, providing robust high-level representations for subsequent detection modules.
In the proposed architecture, CSIF with $n=2$ CSIB blocks operates at the $P_{5}$ fusion level. The hierarchical design—with partial-channel attention preventing detail erosion, CSIB providing stable attention-FFN cycles through post-normalization, and ASRM enabling smooth normalization transition with progressive semantic regulation—collectively achieves the goal of global context modeling without sacrificing the fine-grained discrimination essential for small-object detection. The integration of CSIF at this fusion stage addresses the fundamental limitation of concatenation-based multimodal fusion: while concatenation preserves all information from both modalities, it provides no mechanism for establishing correspondence relationships between RGB-derived and depth-derived feature channels. The global attention computation within CSIF enables direct interaction between features originating from different modalities, learning which RGB patterns should attend to which depth patterns for optimal defect representation. This cross-modal attention emerges implicitly through end-to-end training, as the detection loss gradient encourages attention patterns that support accurate classification and localization of transmission-line defects. The refined $P_{5}$ features subsequently pass through SRM for dictionary-based enhancement, ensuring that both global context coherence from CSIF and reconstructability-based reinforcement from SRM contribute to the semantic representations that propagate through the detection head.

### v3

@@ 2.1.2 Fusion Layers: Selective Cross-Modal Integration at $P_{4}$ and $P_{5}$
Two observations motivate excluding depth from $P_{3}$. First, depth features at this resolution carry the highest sensor noise and spatial misalignment; injecting them into the feature map most critical for small-defect localization would introduce more interference than useful geometric cues. Second, depth primarily contributes structural context—surface discontinuities and inter-component spatial layout—that is adequately captured at $P_{4}$ and $P_{5}$, where feature granularity better matches the sensor’s effective resolution.
At $P_{4}$, the SRM-processed branch features (C=128C{=}128 each) are concatenated along the channel dimension, yielding a 256-channel tensor. A C3k2 block re-encodes this tensor, and a subsequent SRM instance (K=256K{=}256, $\alpha{=}0.8$) refines the fused representation, reinforcing cross-modally consistent patterns while attenuating modality-private residuals.
[…]
2.1.3 Detection Head
The detection head follows the standard YOLO 11 design without modification. In the FPN pathway, fused $P_{5}$ features are upsampled $2\times$ (nearest-neighbor) and concatenated with fused $P_{4}$, then upsampled again and concatenated with RGB-only $P_{3}$; a C3k2 block refines the concatenated features at each level. The PAN pathway reverses this flow via stride-2 convolutions, propagating high-resolution spatial detail from $P_{3}$ back toward $P_{4}$ and $P_{5}$.
Each scale terminates in a decoupled head with separate classification and regression branches. Bounding boxes are predicted via distribution focal loss (DFL), regressing discretized distributions over left-top-right-bottom offsets relative to anchor points. Positive samples are assigned by the Task-Aligned Assigner, which scores candidates jointly by classification confidence and localization quality. The training loss combines binary cross-entropy for classification, CIoU for box regression, and DFL for distribution regression, all with default YOLO 11 weights. The proposed contributions are thus confined entirely to the feature extraction and fusion stages; the detection head and training procedure remain identical to the baseline.
@@ 2.2 Semantic Recomposition Module
SRM is deployed at six locations: $P_{3}$ and $P_{4}$ within each branch (four instances, K=128K{=}128) and the fused $P_{4}$ and $P_{5}$ levels (two instances, K=256K{=}256). All instances share the same four-stage architecture—encode, refine, normalize, decode—followed by residual injection, differing only in the latent dimension $K$.
Figure 2: Architecture of the Semantic Recomposition Module. The four-stage pipeline consists of: (1) pointwise encoding from $C$ to $K$ channels, (2) $5\times 5$ depthwise refinement without activation, (3) position-wise normalization (PONO) across the $K$ dimension, and (4) pointwise decoding from $K$ back to $C$ channels. The output is a convex combination of the refined and original features controlled by the mixing coefficient $\alpha$.
SRM is designed to suppress modality-specific noise without eroding the fine-grained details on which small-defect localization depends. In transmission-line imagery, background textures (conductor strand patterns, fitting reflections), structural edges (conductor–sky boundaries), and depth sensor artifacts (holes, edge bleeding, quantization noise) all produce strong feature responses that inflate false-positive rates at the very scales most relevant to small targets. Channel and spatial attention can reweight existing activations but cannot impose structural constraints on what constitutes valid content. SRM takes a different approach: it compresses features into a learned low-dimensional subspace, refines them within that subspace, and reinjects the result through controlled residual mixing.
@@
Given an input feature map $\mathbf{F}\in\mathbb{R}^{B\times C\times H\times W}$, the pipeline proceeds as follows (Fig. 2). The encoding stage projects $\mathbf{F}$ into a $K$-dimensional latent space via a $1\times 1$ pointwise convolution without BN or activation:
$\mathbf{E}=\mathcal{P}_{\mathrm{enc}}(\mathbf{F}),\quad\mathcal{P}_{\mathrm{enc}}:\mathbb{R}^{C}\rightarrow\mathbb{R}^{K},\quad K<C.$ (1)
The reduction from $C$ to $K$ imposes an information bottleneck that retains only components admitting a stable low-dimensional representation, discarding incidental variation.
The refinement stage aggregates local spatial context within the latent space through a $5\times 5$ depthwise convolution applied independently to each of the $K$ channels:
$\mathbf{R}=\mathcal{D}_{5\times 5}(\mathbf{E}).$ (2)
No activation function follows this convolution: keeping the refinement linear prevents nonlinearity from coupling noise across latent channels. The $5\times 5$ kernel captures local relationships between adjacent structural elements while maintaining parameter efficiency through the depthwise formulation.
@@
The normalization stage applies position-wise standardization across the $K$ dimension at each spatial location:
@@
where $\epsilon=10^{-5}$. Unlike batch or layer normalization, which pool statistics over spatial or batch dimensions, position-wise normalization computes statistics at each $(h,w)$ location independently. Transmission-line images exhibit pronounced spatial heterogeneity—sky, metallic towers, and insulator assemblies produce very different activation patterns—and sharing statistics across these regions would suppress informative variation. A consequence of this design is that absolute activation magnitudes are erased at each position; only the relative distribution across latent channels survives. This reduces sensitivity to the systematic scale difference between RGB and depth activations at corresponding locations, but also discards magnitude information that may be relevant to confidence estimation. The residual mixing in the final stage partially compensates by preserving a fraction of the original features.
The decoding stage maps the normalized representation back to $C$ channels via a $1\times 1$ pointwise convolution (without BN or activation):
@@
The output is a convex combination of the refined and original features:
@@
where $\alpha=0.8$ is a fixed scalar. Weighting the refined branch heavily maximizes noise suppression, while the $(1{-}\alpha)$ residual preserves fine-grained spatial detail that the bottleneck cannot fully reconstruct—a balance between purification strength and localization fidelity.
All six instances share the same architecture but serve distinct functional roles via the choice of $K$. At the branch level (K=128K{=}128), the compact bottleneck filters modality-private noise—illumination artifacts in RGB, sensor holes and quantization errors in depth—by compressing features before normalization, producing per-modality representations with narrower inter-modal distribution gaps. At the fusion level (K=256K{=}256), the expanded bottleneck operates on concatenated cross-modal features; position-wise normalization then acts on the joint RGB-depth latent space, attenuating channels whose activations are inconsistent across modalities while reinforcing those that exhibit coherent patterns. In both cases, $K$ is set to roughly half the input channel dimension, balancing compression strength against representational capacity.
@@ 2.3 Contextual Semantic Integration Framework
Because SRM relies on local convolutions ($5{\times}5$ depthwise kernel) and per-position normalization, it has no mechanism for capturing long-range spatial dependencies. At the $P_{5}$ fusion level, however, defect discrimination often hinges on global structural context: a damaged insulator is easier to identify when the network can reference the regular arrangement of intact neighbors, and a bird nest is distinguished from legitimate equipment by its inconsistency with the expected component layout. To provide this capability, CMAFNet inserts CSIF at the $P_{5}$ fusion stage, between C3k2 and the terminal SRM.
CSIF is placed exclusively at $P_{5}$ for two reasons. Global attention is most effective where features encode abstract relational concepts rather than fine-grained spatial patterns, and the quadratic cost of self-attention ($\mathcal{O}(H^{2}W^{2})$) is tractable only at the lowest resolution ($20{\times}20$ at the nano scale); deploying CSIF at $P_{3}$ or $P_{4}$ would multiply FLOPs by $16\times$ or $4\times$ with no commensurate benefit.
Figure 3: Architecture of the Contextual Semantic Integration Framework. A convolutional stem expands channel dimensions; the resulting feature map is split along the channel axis into an attention portion and a bypass portion. The attention portion is refined by $n$ stacked CSIBs, then concatenated with the bypass and projected to the output dimension.
As illustrated in Fig. 3, CSIF processes an input $\mathbf{X}\in\mathbb{R}^{B\times C_{\mathrm{in}}\times H\times W}$ through a project–refine–project pipeline. A $1{\times}1$ convolutional stem with BN and SiLU expands the channel dimension by a factor of e=2e{=}2, producing a $2C_{\mathrm{in}}$-channel tensor that is then split into two equal halves along the channel axis. One half—the attention portion—passes through a sequence of $n$ Contextual Semantic Interaction Blocks (CSIBs); the other—the bypass portion—is forwarded without modification. After $n$ iterations the two halves are concatenated, and a $1{\times}1$ convolution with BN and SiLU maps the result to the target dimension $C_{\mathrm{out}}$. Restricting attention to half the channels has two benefits. It confines global interaction to a subspace of the representation, mitigating the over-smoothing that full-channel attention would inflict on small-object features, and it halves the computational cost of the attention operation.
@@
Figure 4: Structure of the Contextual Semantic Interaction Block. Each block applies multi-head self-attention followed by ASRM, then a feed-forward network followed by ASRM. Residual connections are present at both stages.
Each CSIB (Fig. 4) implements a standard attention–FFN cycle with ASRM replacing conventional layer normalization in a post-normalization arrangement. The first phase computes multi-head self-attention:
@@
with $h=\max(1,\lfloor C/64\rfloor)$ heads and per-head dimension $d_{k}=C/h$. A residual connection yields $\mathbf{U}=\mathbf{H}+\mathrm{Attn}(\mathbf{H})$, followed by ASRM: $\hat{\mathbf{U}}=\mathrm{ASRM}(\mathbf{U})$. The second phase applies a two-layer feed-forward network:
$\mathrm{FFN}(\hat{\mathbf{U}})=\mathbf{W}_{2}\,\sigma(\mathbf{W}_{1}\hat{\mathbf{U}}+\mathbf{b}_{1})+\mathbf{b}_{2},$ (9)
where $\sigma$ is SiLU and the hidden dimension is $2C$. A second residual connection and ASRM produce the block output: $)\mathbf{H}^{(i)}=\mathrm{ASRM}(\hat{\mathbf{U}}+\mathrm{FFN}(\hat{\mathbf{U}}$)). Post-normalization is chosen over pre-normalization because the concatenated RGB-depth features carry persistent distribution heterogeneity; normalizing after the residual addition corrects this shift before the next transformation can amplify it.
ASRM is designed to regulate the distribution heterogeneity of cross-modal fused features more flexibly than standard layer normalization. Given $\mathbf{x}\in\mathbb{R}^{B\times C\times H\times W}$, it first computes a gated interpolation between the layer-normalized and original representations:
where $\mathrm{LN}(\cdot)$ is channel-wise layer normalization (statistics computed over $C$ at each spatial position), $\boldsymbol{\gamma},\boldsymbol{\gamma}_{x}\in\mathbb{R}^{C\times 1\times 1}$ are learnable per-channel scales initialized to $\boldsymbol{\gamma}{=}0.01$ and $\boldsymbol{\gamma}_{x}{=}1$, and $\odot$ denotes the Hadamard product. With this initialization ASRM approximates an identity mapping at the outset, avoiding premature interference with the still-unstable attention weights; as training proceeds, the gate gradually shifts toward the normalized branch, effecting a smooth transition from weak to strong regulation.
A constrained adaptation mapping $\mathcal{F}$ then computes a low-rank residual correction on the gated features:
@@
where $\mathcal{P}_{\mathrm{down}}{:}\;\mathbb{R}^{C}{\to}\mathbb{R}^{r}$ is a $1{\times}1$ convolution reducing channels to r=64r{=}64, $\mathrm{DW}_{3,5,7}$ denotes three parallel depthwise convolutions with kernel sizes $3{\times}3$, $5{\times}5$, and $7{\times}7$ whose outputs are summed, $\sigma$ is GELU, Dropout operates with probability 0.05, and $\mathcal{P}_{\mathrm{up}}{:}\;\mathbb{R}^{r}{\to}\mathbb{R}^{C}$ restores the original channel dimension. The final output is:
@@
Because $r{=}64\ll C$, the residual formulation confines the adaptation to a low-rank subspace, ensuring that $\mathcal{F}$ acts as a correction rather than a full feature rewrite. The multi-scale depthwise convolutions supply spatially adaptive context at three receptive-field sizes without the quadratic cost of a second attention layer. This combination distinguishes ASRM from both standard gated normalization—which lacks a spatial adaptation path—and adapter-style modules—which lack the gated normalization front-end that governs the identity-to-regulation transition.
CSIF uses n=2n{=}2 CSIB blocks at $P_{5}$, balancing modeling capacity against the over-smoothing that additional blocks would introduce at $20{\times}20$ resolution. Through global attention, the module enables direct interaction between RGB-derived and depth-derived features across all spatial positions, letting the network learn cross-modal correspondences beneficial to defect detection. The refined $P_{5}$ features then pass through the terminal SRM for a final round of bottleneck-based distribution normalization before entering the detection head.